Can someone please just explain me how to resolve this? Thank you.
The charge to rent a trailer is $30 for up to 2 hours plus $9 per additional hour or portion on an hour. Find the cost to rent a trailer for 3.7 hours, 4 hours and 8.6 hours.
Then graph all ordered parts.
What is the cost to rent a trailer for: a) 3.7 hours? b) 4 hours?

Answer by rothauserc(4718)

3.7 = 2 * 30 + 2 * 9 = $78
4.0 = 2 * 30 + 2 * 9 = $78
8.6 = 2 * 30 + 7 * 9 = $123
:
(3.7, 78), (4, 78), (9, 123)
:
a) cost = 2 * 30 + 2 * 9 = $78
:
b) cost = 2 * 30 + 2 * 9 = $78